OpenStudy (anonymous):

The label on a bottle reads "MOH", and you set out to find the identity of the metal (M). Assuming that this is an alkali metal hydroxide, you weigh out 1.67g of the solid. It takes 0.0418L of 1.00M HCl to titrate the mtal hydroxide. What is the identity of the solid?

OpenStudy (jfraser):

If you assume that "M" is a group 1 metal, then the balanced reaction will be \[MOH + HCl \rightarrow MCl + H{_2}O\] you know t he volume and concentration of acid added, so you know moles of acid added. You know the reaction is 1:1, so the moles of acid will be the same as the moles of metal "M"OH used. You know the mass of metal hydroxide used, and you know the moles used, so divide mass/mols to get molar mass and identify the compound.\[vol Acid * (\frac{mol acid}{vol acid})* (\frac{mol MOH}{mol acid}) = moles MOH used\] \[MM = \frac{mass MOH used}{moles MOH calculated}\]